OpenStudy (wolf1728):

Trying to solve for 'n'. (r*pv)/pmt = 1-((1+r)^-n) if r=.01; pv=4,495.50;pmt=100;n=60 then (.01*4,495.50)/100 =.44955 and 1-((1+r)^-n) equals 1 -(1.01)^-60 equals 0.44955 I am trying to solve the equation for 'n' which requires taking logs of both sides. The log of (r*pv)/pmt is easy enough to find but trying to find the log of the right side of the equation 1-((1+r)^-n) is a lot more difficult. Ordinarily the log should be -n* (1 - (1+r)). However, the "1" the "-(1+r)" and the "-n" are presenting problems. (Negative numbers present severe problems when dealing with logarithms. Anyone have any idea how to solve the equation for "n"?

OpenStudy (mathstudent55):

\(\dfrac{r \times pv}{pmt} = 1-(1+r)^{-n}\) Add the quantity with n in the exponent to both sides and subtract the left fraction from both sides to have all terms with n on the left side: \((1+r)^{-n}= 1-\dfrac{r \times pv}{pmt} \) 1 = pmt/pmt, and add the fractions onthe right side. \((1+r)^{-n}= \dfrac{pmt - r \times pv}{pmt} \) Take logs of both sides: \(\log [(1+r)^{-n}]= \log[\dfrac{pmt - r \times pv}{pmt}] \) \(-n\log (1+r)= \log[\dfrac{pmt - r \times pv}{pmt}] \) \(-n= \dfrac{\log[\dfrac{pmt - r \times pv}{pmt}]}{ \log (1+r) } \) \(-n= \dfrac{\log( pmt - r \times pv) - \log pmt}{ \log (1+r) } \) \(n= \dfrac{\log pmt -\log( pmt - r \times pv) }{ \log (1+r) } \) Using your values, and solving for n: \(n= \dfrac{\log 100 -\log( 100 - 0.01 \times 4495.5) }{ \log (1+0.01) } \) \(n= \dfrac{2 -\log 55.045 }{ \log (1.01) } \) \(n = 59.99992989\) n comes out very close to 60.

OpenStudy (wolf1728):

Wow thanks. Yes, that was a little bit "beyond" me.

OpenStudy (mathstudent55):

yw